# Modified Marrone-Treanor dissociation model: formulation and verification for diatom/atom mixtures [^1] Portions of this work were presented as papers 2019-0789 and 2020-2191 at AIAA SciTech forums held in San Diego, CA. and Orlando, FL respectively

### v1

@@ 3.3 Role of Vibration
[…]
For a given internal energy, molecules that are more vibrationally excited are found to be more likely to dissociate; this conclusion is consistent with our findings using the support factor. For example, we identify two sample rovibrational levels with similar internal energy: one with small $\varepsilon_{v}$ and the other with large $\varepsilon_{v}$. The state which has its internal energy in rotation is approximately 8.5 times less likely to dissociate than the state which has its internal energy in vibration. The effective diatomic potentials for these two states are shown in Fig. 3. Although neither of these molecules is quasibound, the molecule with mostly rotational energy requires more energy to dissociate due to the centrifugal barrier, assuming its angular momentum remains constant. We quantify this by defining the remainder energy, $\varepsilon_{\mathrm{rem}}$, which is the difference between the centrifugal barrier height $V_{\mathrm{D,eff}}$ and the internal energy $\varepsilon_{\mathrm{int}}$,
@@
in which function $\max[.]$ refers to the maximum in the dissociation barrier; see Figure 3.
Several analyses in the literature have used the barrier height in theoretical arguments, often as an effective dissociation energy [38, 39, 10]; this interpretation is similar. Here, we directly evaluate its validity as a predictor for dissociation using the QCT data. The probability of dissociation compared to the remainder energy is shown in Fig. 4. We find that the barrier energy is a relatively good predictor of dissociation; the scatter in probability is approximately a factor of 2, a substantial portion of which is due to statistical uncertainty. Similar observations are made for the other single dissociation [^6] Single dissociation refers to quasiclassical trajectories of type $\mathrm{AB}+\mathrm{CD}\rightarrow\mathrm{A}+\mathrm{B}+\mathrm{CD}$, as opposed to those involving simultaneous exchange-dissociation: $\mathrm{AB}+\mathrm{CD}\rightarrow\mathrm{A}+\mathrm{C}+\mathrm{BD}$, or double-dissociation: $\mathrm{AB}+\mathrm{CD}\rightarrow\mathrm{A}+\mathrm{B}+\mathrm{C}+\mathrm{D}$. reactions we have considered.
Therefore, the remainder energy is a good predictor for single dissociation. This result also indicates an important physical mechanism: the reason why rotational energy is less effective at promoting dissociation is because rotational energy also increases the centrifugal barrier. Thus, adding one unit of rotational energy reduces $\varepsilon_{\mathrm{rem}}$ by an amount less than the energy added. This reduction does not occur for vibrational energy.
@@ 4 Fitting to QCT Data
In this section, we describe the methodology and results for fitting aggregate quantities of interest from QCT to relatively simple analytic forms. When possible, functional forms that exist in the literature are used. Fits are considered for both the dissociation rate and the change in vibrational energy per dissociation for the equilibrium and nonequilibrium test sets. Each dissociation reaction with a specific collision partner is considered separately.
@@ 4.1 Equilibrium Dissociation Rate Coefficient
We first fit dissociation rate data for the equilibrium test set ($T=T_{v}$) to the modified Arrhenius form, Eq. (11). We write the equation in terms of inverse temperature and the natural logarithm of the dissociation rate coefficient,
$\ln k_{d}^{\mathrm{arr}}=\ln C-n\ln\frac{1}{T}-\frac{T_{D,s}}{T}.$ (22)
We choose to use this form for several reasons. First, the equation is now linear in the coefficients we wish to determine: $\ln C$, $-n$, and $-T_{D,s}$. We can therefore use a standard least squares estimator to determine the coefficients (e.g., Riley et al. [40] [p. 1271–1277]). Second, a least squares minimization in $\ln k_{d}^{\mathrm{arr}}$ corresponds to minimizing the relative error in $k_{d}^{\mathrm{arr}}$, which is desirable. For the temperature ranges of interest, $k_{d}^{\mathrm{arr}}$ varies by many orders of magnitude, so minimizing the absolute error would essentially neglect relative error in the small but important $k_{d}^{\mathrm{arr}}$ values at low temperatures.
Results for all dissociation reactions we have considered are shown in Table 2. The data are found to be well described by the modified Arrhenius form; the goodness-of-fit metric $R^{2}$ (computed in $\ln k_{d}^{\mathrm{arr}}$ space) is better than $0.9998$ for all results shown. For all effective dissociation rate coefficients, the maximum deviation is less than 5%, and it is often better than the estimated statistical QCT error for low-temperature conditions. The fitted values of $n$ are typically between 00 and $-1.5$, which is consistent with theoretical results (e.g. Ref. [41]) and data available in the literature (e.g., Ref. [42]). Note that the fitted dissociation temperature is within 5% of the dissociation temperature based on the diatomic PES’s dissociation energy. In the experimental literature, it is common to fit Arrhenius coefficients with a fixed $T_{D,s}$. We have done this with accepted values of $T_{D,s}$, and find that the fits are only slightly changed relative to those shown in Table 2.
@@ 4.2 Nonequilibrium Dissociation Rate Coefficient
Next, we consider the dissociation rate coefficient for the nonequilibrium QCT dataset. The rate coefficient at thermal nonequilibrium is often considered relative to the one at thermal equilibrium by defining and fitting a nonequilibrium correction factor, Eq. (12). The dissociation rate coefficients at equilibrium are defined by the Arrhenius form, with values taken from the previous section’s fits. Similar to the equilibrium rate coefficient, we fit in terms of the natural logarithm (i.e. $\ln Z$). A variety of nonequilibrium correction factors are defined in the literature; the book chapter by Losev et al. [43] provides a good summary. Here, we focus on the Marrone-Treanor model, because it was found to be consistent with the QCT data. Prior work [44] compared the Losev $\beta$ model [7], the Park $T-T_{v}$ model [45, 46], and the more recent Luo et al. [10] model.
@@
$\ln Z(T,T_{v})=\ln Q_{v}(T)+\ln Q_{v}(T_{F})-\ln Q_{v}(T_{v})-\ln Q_{v}(-U)$ (23)
@@
The results for effective oxygen dissociation with all collision partners are shown in Fig. 5; we have chosen to show oxygen dissociation for brevity, but nitrogen dissociation has been analyzed in the same fashion. The Marrone-Treanor model is found to fit the data well: the maximum relative deviation is within 5% for nitrogen dissociation with partners $\mathrm{N_{2}}$ and $\mathrm{N}$ and oxygen dissociation with partners $\mathrm{O_{2}}$ and $\mathrm{O}$, and within 7.5% for both dissociation reactions of $\mathrm{N_{2}+O_{2}}$ pairs.
@@
Figure 5: Nonequilibrium correction factor $Z$ and average change in vibrational energy per dissociation $\langle\varepsilon_{v,s}\rangle_{d}$ for effective oxygen dissociation in the nonequilibrium test set.
@@ 4.3 Nonequilibrium Vibrational Energy Change per Dissociation
The Marrone-Treanor model yields for the vibrational energy removed per dissociation:
@@
$\langle\varepsilon_{v,s}\rangle_{d}(T,T_{v})=\frac{1}{Q(T_{F})}\sum_{v}\exp\left(-\frac{\varepsilon_{s}(v)}{\mathrm{k_{B}}T_{F}}\right),$ (24)
which requires one to perform a Boltzmann-weighted sum spanning all the diatom’s vibrational energy levels $\varepsilon_{s}(v)$ at pseudo-temperature $T_{F}(T,T_{v})$. However, this costly calculation can be avoided by approximating Eq. (24) with Eq. (16), the form originally given by Knab et al. [9].
@@
As before, the Marrone-Treanor model is fitted for each dissociation reaction, with the fitted dissociation temperature. The characteristic vibrational temperature $\theta_{v,s}$ is defined using the corresponding diatomic PES and the dissociation energy is described using the dissociation temperature, $T_{D,s}=D_{0,s}/\mathrm{k_{B}}$. The results for oxygen dissociation with all collision partners are shown in Fig. 5; the results for nitrogen dissociation are similar. The Marrone-Treanor model is found to accurately describe the data: the maximum error is $0.13\,\mathrm{eV}$ for oxygen dissociation and $0.49\,\mathrm{eV}$ for nitrogen dissociation. The best-fit model parameter $U$ is also found to be similar when fitting to rate data and when fitting to change in energy data (within 7% for oxygen dissociation and 15% for nitrogen dissociation). For all cases, the fitted model parameter is between $\frac{1}{3.5}$ and $\frac{1}{2}$ of the dissociation temperature.
4.4 Marrone-Treanor Pseudo-Temperature
Based on these results, we propose a modification to the Marrone-Treanor preferential dissociation model that yields a fit to both quantities of interest, in both the equilibrium and nonequilibrium test set, for all effective dissociation reactions that we have considered. Both the reaction rate coefficient and the change in vibrational energy per dissociation are well described by a single value of $U$, when $T$ is fixed. By construction, this value of $U$ also fits $\langle\varepsilon_{v,s}\rangle_{d}$ at thermal equilibrium for the single condition that matches the value of $T$ used for the nonequilibrium test set. Therefore, the value of $U$ that describes the nonequilibrium behavior for a given $T$ may be determined by matching only the $\langle\varepsilon_{v,s}\rangle_{d}$-value at the corresponding equilibrium condition.

### v2

@@ 3.3 Role of vibration
In Secs. 3.1 and 3.2 we presented aggregate QCT-derived data for vibrational energy change and support factor for all six dissociation reactions in Table 2 over a wide range of temperatures. Perhaps unsurprisingly, the overall message of this analysis is that the vibrational energy of the dissociating diatom has the strongest effect on dissociation probability. In this subsection, we attempt to explain why vibrational energy has a stronger effect on dissociation than rotational energy. For this discussion we use the $\mathrm{O_{2}+O\rightarrow 3O}$ reaction, averaged over all ground-electronic-state interactions at $T=T_{\mathrm{v}}=10\,000\,\mathrm{K}$.
Figure 3.33 of Ref. [25] shows a contour plot of dissociation probability as a function of rovibrational level $(v,j)$ for this reaction. The molecules most likely to dissociate (orange to red shaded area) populate levels with either very large $v$ (bottom right corner), very large $j$ (top left corner) or $(v,j)$ combinations running along the upper-right edge of the plot. These molecules all have in common that they possess internal (i.e. vibrational+rotational) energies close to, or above the dissociation threshold $D_{0}$ for $\mathrm{O_{2}}$. Furthermore, as shown in greater detail by Figs. 3.34(a) and (b) of Ref. [25], we find that the probability of dissociation from each rovibrational level is not fully predicted by either vibrational or rotational energy alone.
For a given internal energy, we find molecules that are more vibrationally excited to be more likely to dissociate; this conclusion is consistent with our findings using the support factor. For example, we identify two sample rovibrational levels with similar internal energy, one with small $\varepsilon_{\mathrm{vib}}$ and the other with large $\varepsilon_{\mathrm{vib}}$ (see Fig. 3.35(b) in Ref. [25]). The state which has most of its internal energy in rotation is approximately $8.5$ times less likely to dissociate than the state which has its internal energy in vibration. We show the effective diatomic potentials for these two states in Fig. 4. Although neither of these molecules is quasibound, the molecule in Fig. 4(b), with mostly rotational energy requires more energy to dissociate due to the centrifugal barrier, assuming its angular momentum remains constant. We quantify this by defining the remainder energy, $\varepsilon_{\mathrm{rem}}$, which is the difference between the centrifugal barrier height and the internal energy $\varepsilon_{\mathrm{int}}=\varepsilon_{\mathrm{vib}}+\varepsilon_{\mathrm{rot}}$,
@@
in which function $\max[V_{\mathrm{D,eff}}(j)]$ refers to the local maximum of the effective diatomic potential curve; see Figure 4.
Several analyses in the literature have used the barrier height in theoretical arguments, often as an effective dissociation energy [46, 47, 10] and our interpretation is similar. Here we directly evaluate its validity as a predictor for dissociation using the QCT data. We show the probability of dissociation compared to the remainder energy in Fig. 5. We find that the remainder energy is a relatively good predictor of dissociation; the scatter in probability is approximately a factor of 2, a substantial portion of which is due to statistical uncertainty. We made similar observations for the other dissociation reactions we have considered. Thus, the remainder energy appears to be a good predictor for single dissociation (see footnote11footnotemark: 1 in Sec. 2.2). This result also indicates an important physical mechanism: the reason why rotational energy is less effective at promoting dissociation is because rotational energy also increases the centrifugal barrier. Thus, adding one unit of rotational energy reduces $\varepsilon_{\mathrm{rem}}$ by an amount less than the energy added. This reduction does not occur for vibrational energy.
Admittedly, the preceding explanation has its limitations. Whether any specific molecule splits apart will not only be a function of its effective diatomic potential curve and associated remainder energy, but will depend in a non-trivial manner on the precise collision dynamics, i.e. the collision pair’s relative translational energy, the spatial orientation of the collision partner’s incoming trajectory relative to the molecule, as well as the timing of closest approach. Even after individual trajectories have been aggregated to average out these dynamic effects, Fig. 5 represents only a single equilibrium temperature condition. When $T$ is changed, it shifts the relative translational energy distribution of colliding pairs. This will undoubtedly change the aforementioned dissociation probability ratio of $8.5$ between the two chosen levels, as well as affect the overall $P_{d}$ vs. $\varepsilon_{\mathrm{rem}}$ curve. Thus, the height of the centrifugal barrier is likely not the only factor for explaining dissociation probability. Regardless of the precise reasons, the aggregate QCT results presented in this section confirm the strong vibrational bias in dissociation reactions. This ultimately justifies the major modeling choices presented in Sec. 2.2 for the MMT model.
@@ 4 Fitting to QCT data
In this section, we continue summarizing parts of chapter 3 (in particular Sec. 3.6) of Ref. [25]. We describe the methodology and results for fitting aggregate quantities of interest from QCT to relatively simple analytic forms. When possible, we use functional forms that exist in the literature. We consider fits for both the dissociation rate and the change in vibrational energy per dissociation for the equilibrium and nonequilibrium test sets. But, as we will show, we can extract the model parameters appearing in Table 2 from just the equilibrium test sets for dissociation rate coefficient and energy removed by dissociation. We consider each dissociation reaction with a specific collision partner separately.
@@ 4.1 Equilibrium dissociation rate coefficient
We first fit dissociation rate data for the equilibrium test set ($T\!=\!T_{\mathrm{v}}$) to the modified Arrhenius form of Eq. (11). We write the equation in terms of inverse temperature and the natural logarithm of the dissociation rate coefficient,
$\ln k^{\mathrm{Arr}}=\ln C-n\ln\frac{1}{T}-\frac{T_{\mathrm{D},s}}{T}.$ (22)
We choose to use this form for several reasons. First, the equation is now linear in the coefficients we wish to determine $\ln C$, n-n, and $T_{\mathrm{D},s}$. We can therefore use a standard least squares estimator to determine the coefficients (e.g., Riley et al. [48] [p. 1271–1277]). Second, a least squares minimization in $\ln k^{\mathrm{Arr}}$ corresponds to minimizing the relative error in $k^{\mathrm{Arr}}$, which is desirable. For the temperature ranges of interest, $k^{\mathrm{Arr}}$ varies by many orders of magnitude, so minimizing the absolute error would essentially neglect relative error in the small but important $k^{\mathrm{Arr}}$ values at low temperatures.
We show results for all dissociation reactions we have considered in Table 2. We find the data to be well described by the modified Arrhenius form; the goodness-of-fit metric $R^{2}$ (computed in $\ln k^{\mathrm{Arr}}$ space) is better than $0.9998$ for all results shown. For all effective dissociation rate coefficients, the maximum deviation is less than 5%, and it is often better than the estimated statistical QCT error for low-temperature conditions. The fitted values of $n$ are typically between 0 and 1.5-1.5, which is consistent with theoretical results (e.g. Ref. [49]) and data available in the literature (e.g., Ref. [50]). Note that the fitted dissociation temperature in Table 2(a) is within 5% of the dissociation temperature based on the diatomic PES’s dissociation energy. In the experimental literature, it is common to fit Arrhenius coefficients with a fixed $T_{\mathrm{D},s}$. We have done this with accepted values of $T_{\mathrm{D},s}$ in Table 2(b), and find that the fits are only slightly changed relative to those shown in Table 2(a). A more detailed listing containing all modified Arrhenius fit parameters obtained with and without a fixed $T_{\mathrm{D},s}$ can be found in Tables 3.1 and 3.2 of Ref. [25].
@@ 4.2 Nonequilibrium dissociation rate coefficient
Next, we consider the dissociation rate coefficient for the nonequilibrium QCT dataset. The rate coefficient at thermal nonequilibrium is often considered relative to the one at thermal equilibrium by defining and fitting a nonequilibrium correction factor, Eq. (12). We define the dissociation rate coefficients at equilibrium by the Arrhenius form, with values taken from the previous section’s fits. Similar to the equilibrium rate coefficient, we fit in terms of the natural logarithm (i.e. $\ln Z$). A variety of nonequilibrium correction factors are defined in the literature; the book chapter by Losev et al. [51] provides a good summary. Here we focus on the Marrone-Treanor model, because we found it to be consistent with the QCT data. Prior work [25] compared the Losev $\beta$ model [7], the Park $T-T_{\mathrm{v}}$ model [52, 53], and the more recent Luo et al. [10] model.
@@
$\ln Z(T,T_{\mathrm{v}})=\ln\tilde{Q}_{\mathrm{v},s}(T)+\ln\tilde{Q}_{\mathrm{v},s}(T_{F})-\ln\tilde{Q}_{\mathrm{v},s}(T_{\mathrm{v}})-\ln\tilde{Q}_{\mathrm{v},s}(-U)$ (23)
@@
In Fig. 6(a) we show the curve fit to $Z(T,T_{\mathrm{v}})$ for effective oxygen dissociation with collision partner $\mathrm{N_{2}}$ at $T=10\,000\,\mathrm{K}$. The best fit to the QCT data is achieved for $U=20\,836\,\mathrm{K}$. We have chosen to show just this one reaction at a single trans-rotational temperature for brevity, but equivalent plots for $Z$ in $\mathrm{O_{2}}$ dissociation with partners $\mathrm{O_{2}}$ and $\mathrm{O}$ can be found in Fig. 3.38(a) and 3.38(b) of Ref. [25]. We have further analyzed nitrogen dissociation with the collision partners listed in Table 2 in the same fashion (not shown) and found the Marrone-Treanor model to fit the data well; the maximum relative deviation is within 5% for nitrogen dissociation with partners $\mathrm{N_{2}}$ and $\mathrm{N}$ and oxygen dissociation with partners $\mathrm{O_{2}}$ and $\mathrm{O}$, and within 7.5% for both dissociation reactions of $\mathrm{N_{2}+O_{2}}$ pairs.
@@
Figure 6: Nonequilibrium factor $Z$ and average change in vibrational energy per dissociation $\langle\varepsilon_{\mathrm{v},s}\rangle_{\mathrm{diss}}$ for effective oxygen dissociation in nonequilibrium test set.
@@ 4.3 Nonequilibrium vibrational energy change per dissociation
The Marrone-Treanor model yields for the vibrational energy removed per dissociation (see Eq. (8) in Ref.[5]):
@@
$\langle\varepsilon_{\mathrm{v},s}\rangle_{\mathrm{diss}}\left(T_{F}\right)=\frac{1}{Q_{\mathrm{v},s}(T_{F})}\sum_{v=0}^{n_{v,s}-1}\varepsilon_{s}(v)\exp\left(-\frac{\varepsilon_{s}(v)}{\mathrm{k_{B}}T_{F}}\right).$ (24)
The ground-electronic-state $\mathrm{N_{2}}$ diatomic potential of the Minnesota PESs accommodates nearly 60 discrete vibrational levels, whereas for $\mathrm{O_{2}}$ it is about 45 (see supplemental information of Ref. [37] for detailed listings). The exact formulation given by Eq. (24) requires one to perform two separate Boltzmann-weighted sums spanning all the diatom’s vibrational energy levels $\varepsilon_{s}(v)$ at pseudotemperature $T_{F}(T,T_{\mathrm{v}})$, one for the numerator and another one for the vibrational partition function:
$Q_{\mathrm{v},s}(T_{F})=\sum_{\mathrm{v}=0}^{n_{v,s}-1}\exp\left(-\frac{\varepsilon_{s}(v)}{\mathrm{k_{B}}T_{F}}\right)$ (25)
in the denominator. However, Knab et al. [9] avoid this costly calculation by approximating the exact state sums with simpler expressions derived from sums over simple harmonic oscillator (SHO) levels $\varepsilon_{s}(v)]^{\mathrm{SHO}}=v\,\mathrm{k_{B}}\Theta_{\mathrm{v},s}$ truncated below the dissociation threshold $D_{0,s}$. This is how they arrive at the approximate expression for the partition function given by Eq. (13) and are able to approximate Eq. (24) with Eq. (16), the form originally given as Eq. (39) in Knab et al. [9].
In Fig. 7(a) we compare predictions for $\langle\varepsilon_{\mathrm{v},s}\rangle_{\mathrm{diss}}\,(T,T_{\mathrm{v}})$ for reaction $\mathrm{2N_{2}\rightarrow 2N+N_{2}}$ from the exact expression, i.e. Eq.(24), against Knab’s approximate formula, i.e. Eq. (16). We conduct sweeps over $T_{\mathrm{v}}\leq T$ for three distinct values of $T$. Knab’s formula (dotted lines) consistently under-predicts the values obtained with the exact expression (solid lines). At all three translational temperatures shown, the gap between both curves grows with increasing vibrational temperature and becomes largest at thermal equilibrium. But even at $T=T_{\mathrm{v}}=15\,000\,\mathrm{K}$ (black curves), the absolute energy difference only maxes out at about 10% of $D_{0,\mathrm{N_{2}}}$.
In Fig. 7(b) we compare the result of evaluating Eq. (12) for the nonequilibrium factor $Z(T,T_{\mathrm{v}})$ using the exact vibrational partition functions according to Eq. (25), vs. the truncated SHO ones from Eq. (13). At $T_{\mathrm{v}}\ll T$ the approximate formula (dotted lines) predicts slightly larger values than the exact one (continuous lines). This is most clearly visible for $T=15\,000\,\mathrm{K}$ (black lines). However, in absolute terms the error in $Z$ remains small and vanishes at thermal equilibrium.
(a) Vibrational energy loss in dissociation from exact expression [Eq. (24)] vs. Knab’s formula [Eq. (16)].
(b) Nonequilibrium factor [Eq. (12)] computed using exact expressions for $Q_{\mathrm{v,N_{2}}}$ [Eq. (25)] vs. approximate partition functions $\tilde{Q}_{\mathrm{v,N_{2}}}$ [Eq. (13)].
Figure 7: Exact (solid lines) vs. approximate (dotted lines) expressions for vibrational energy loss $\langle\varepsilon_{\mathrm{v,N_{2}}}\rangle_{\mathrm{diss}}(T,T_{\mathrm{v}})$ and nonequilibrium factor $Z(T,T_{\mathrm{v}})$. Parameters for reaction $\mathrm{2N_{2}\rightarrow 2N+N_{2}}$ from Table 2(a).
Despite the discrepancies, switching to the approximate expressions is still worthwhile, as it is extremely useful in reducing the computational cost of the MMT model in large-scale CFD computations, where Eqs. (13) and (16) must be evaluated for every dissociation reaction multiple times per cell and per time step. Furthermore, by applying the SHO approximation in this manner we avoid the need for including detailed listings of $\varepsilon_{s}(v)$ for every dissociating species as part of the two-temperature model. Instead, we merely require their characteristic vibrational and dissociation temperatures.
In hindsight, the preceding discussion justifies fitting $\langle\varepsilon_{\mathrm{v},s}\rangle_{\mathrm{diss}}(T,T_{\mathrm{v}})$ to Eq. (16) for the Marrone-Treanor model. We focus on results generated with the variable dissociation temperatures listed in Table 2(a). An equivalent fit using the fixed $T_{\mathrm{D,s}}$ values from Table 2(b) can be performed, but is not shown here. As mentioned in Sec. 2.2, the characteristic vibrational temperature $\Theta_{\mathrm{v},s}$ is defined using the energy gap between the two lowest-lying vibrational levels derived from the corresponding diatomic potential.
The results for oxygen dissociation with collision partner $\mathrm{N_{2}}$ are shown in Fig. 6(b). QCT data from the nonequilibrium test set are shown as black dots and the best fit to Eq. (16) with $U=21\,399\,\mathrm{K}$ is shown in green. Equivalent plots for $\langle\varepsilon_{\mathrm{v},\mathrm{O_{2}}}\rangle_{\mathrm{diss}}$ in $\mathrm{O_{2}}$ dissociation with partners $\mathrm{O_{2}}$ and $\mathrm{O}$ can be found in Figs. 3.38(d) and 3.38(e) of Ref. [25]. Just as for oxygen, we have analyzed $\langle\varepsilon_{\mathrm{v},\mathrm{N_{2}}}\rangle_{\mathrm{diss}}$ for nitrogen dissociation with the collision partners listed in Table 2 in the same fashion (not shown).
@@
Overall, we found the Marrone-Treanor model to accurately describe the data: the maximum error is $0.13\,\mathrm{eV}$ for oxygen dissociation and $0.49\,\mathrm{eV}$ for nitrogen dissociation. We found the best-fit model parameter $U$ to be similar when fitting to rate data and when fitting to change in energy data (within 7% for oxygen dissociation and 15% for nitrogen dissociation). For all cases, the fitted model parameter is between $\frac{1}{3.5}$ and $\frac{1}{2}$ of the dissociation temperature.
4.4 Marrone-Treanor characteristic probability temperature
In the original derivation of their preferential dissociation model Marrone and Treanor [5] had to rely on shock tube data to infer $U$ (what they called the characteristic probability temperature) essentially by trial and error. Given the large uncertainty in the experimental data, this meant they could only obtain a single, rough approximate value for a given reaction instead of a generally valid functional dependence for $U$. By contrast, with the help of our extensive QCT database we have the ability to calculate $U$ from first principles and may examine its variation over a wide range of nonequilibrium conditions. Based on the results of Secs. 4.2-4.3, we thus propose a modification to the Marrone-Treanor model meant to yield a good fit to both quantities in the equilibrium ($T\!=\!T_{\mathrm{v}}$) and nonequilibrium ($T\!\neq\!T_{\mathrm{v}}$) test sets and for all effective dissociation reactions considered.
As was seen in Fig. 6, the nonequilibrium factor and vibrational energy change per dissociation are each well described by a particular value of $U$ at a fixed value of $T$ and the two lie fairly close to one another. Specifically, for $T=10\,000\,\mathrm{K}$ the value of $U$ that simultaneously fits both quantities well, lies somewhere between $20\,836\,\mathrm{K}$ and $21\,399\,\mathrm{K}$. Similar behavior was observed at other temperatures and for the other reactions studied.
The goal is now to find a simple expression that allows us to estimate $U$ (and by extension $T_{F}$) with good enough accuracy for any $T,T_{\mathrm{v}}$ combination and for any reaction within the entire QCT data set. We face the choice of whether to base our analysis on fitting to data for $Z$, for $\langle\varepsilon_{\mathrm{v},s}\rangle_{\mathrm{diss}}$, or some combination thereof. Because we are primarily interested in minimizing the error in vibrational energy decrease per dissociation, we opt for fitting to this quantity and proceed as follows. By construction the value of $U=21\,399\,\mathrm{K}$ in Fig. 6(b) also fits $\langle\varepsilon_{\mathrm{v},s}\rangle_{\mathrm{diss}}$ at thermal equilibrium, i.e. the single point where $T_{\mathrm{v}}\!=\!T=\!10\,000\,\mathrm{K}$. Thus, for a given temperature $T$ we may find the value of $U$ that describes the entire nonequilibrium behavior by only having to match the value for $\langle\varepsilon_{\mathrm{v},s}\rangle_{\mathrm{diss}}$ from QCT at the corresponding equilibrium condition.